1. A certain amount if nitrous acid is dissolved in water at 25 celcius. At equilibrium, the pH of the solution is 2.33. What is the concentration of H3O+ and NO2-?

2. Calculate the valye if the equilibriym constant at 25 celcisus for the reaction: HNO2+OH- -> H2O + NO2-
Answer #1

1. HNOQ (aq) + H2O(l) = H20+(aq) + Noz (9) Hf HNO2. is dissolved, then [H30*3= [NO27 Now pH=-log [H30*3= 2.33 Or, [H30*7=1023N.B.: As per Chegg guidelines we have to answer first question out of multiple questions. Though we tried to answer the 2nd too, but there is some type mistakes. Without proper question and adequate data we can't answer. But we expect you can solve the rest part. Others reupload it woth proper expression. All the best.